OpenStudy (anonymous):

use the trigonometric ratios for a right triangle to derive a cofunction identity for tan(90degrees-A)

OpenStudy (jack1):

so do you know SOH CAH TOA? in this case TOA ==> Tan (theta) = Opposite/ Adjacent so theta = 90 - A which means its a right angle triangle with 90 degrees for one angle, A degrees for the other angle, and (90-A) degrees for the 3rd angle

OpenStudy (jack1):

 does this make sense???

OpenStudy (anonymous):

that makes sense but how do I derive the cofunction identity using the trig ration\[tantheta=1/\cot\]

OpenStudy (mathmate):

Another hint: \[tan \theta=opp/adj\] \[cot \theta=adj/opp\] and study Jack's diagram again.

OpenStudy (anonymous):

I'm just having trouble applying this, how can I figure any side out from just one angle?